Two 143-cm-long thin glass rods uniformly charged to +123 pC are placed side by side, 13 cm apart. What is the electric field strengths (in N/m) at a distance of 1.8 cm to the right of the rod on the left along the line connecting the midpoints of the two rods?
Answer #1

9- q=123 Pc + + EL 1.8cm E 11.2cm- 2 = 13-118 + 1 13cm - 11.2cm x = q = 123 PC 1043 L E, =) Ez 2 s ZAGNI 2002 Enet 3 Datori P